OpenStudy (anonymous):

∫ 1 between 0 of : t/√(4-t²) dt

ganeshie8 (ganeshie8):

sub u = 4-t^2

OpenStudy (anonymous):

that's what I did but it doesn't work ...

ganeshie8 (ganeshie8):

lets do it again and see...

OpenStudy (anonymous):

the answer is : 0,27 and that's why this doesnt work :(

ganeshie8 (ganeshie8):

u = 4-t^2 du = -2t dt -du/2 = tdt as t->0 u -> 4-0^2 = 4 as t->1 u-> 4-1^2 = 3

ganeshie8 (ganeshie8):

^corrected the typo above

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_1^0 \frac{t}{\sqrt{4-t^2}}dt}\) \(\large \mathbb{\int_3^4 \frac{1}{\sqrt{u}}(-du/2)}\)

ganeshie8 (ganeshie8):

evaluate the integral ?

OpenStudy (anonymous):

that's what I try to do :P

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_3^4 \frac{1}{\sqrt{u}}(-du/2)}\) \(\large \mathbb{\frac{-1}{2}\int_3^4 u^{\frac{-1}{2}}du}\)

ganeshie8 (ganeshie8):

try now :)

OpenStudy (anonymous):

but where did you took your 4/3? Cuz I have the same equation if it where not this :o

OpenStudy (anonymous):

(not good in english, sorry :) )

ganeshie8 (ganeshie8):

i get exactly what u mean lol, ur english is good ok

ganeshie8 (ganeshie8):

u are asking how did i get 3 and 4 in place of 1 and 0 right ?

OpenStudy (anonymous):

Umm kind of yea

ganeshie8 (ganeshie8):

we substituted u = 4-t^2 that means, we have introduced a new 'u' variable in place of 't' so the bounds also will change. when t = 0, u need to find whats the value of u when t = 1, u need to find whats the value of u

ganeshie8 (ganeshie8):

basically when u do substitutions, u need to carry out below steps :- 1) substitute u 2) find du in terms of dt 3) find how bounds will change 4) evaluate the new integral in u's 5) convert back to t's

ganeshie8 (ganeshie8):

if that confuses u, forget it. just see that u are substituting u = 4-t^2 so u need to find the new bounds.

ganeshie8 (ganeshie8):

u = 4-t^2 when t = 0, u = ?

OpenStudy (anonymous):

4`:) haha

ganeshie8 (ganeshie8):

yes, when t = 1, u = ?

OpenStudy (anonymous):

3

ganeshie8 (ganeshie8):

good :)

ganeshie8 (ganeshie8):

so, bounds from 0->1 will change to 4->3

ganeshie8 (ganeshie8):

fine ? are u at peace wid the bounds ha ? :)

OpenStudy (anonymous):

so when you have the new bounds, you do the soustraction?

ganeshie8 (ganeshie8):

u mean, when u substitute, u need to change the bounds accordingly ? then yes :)

ganeshie8 (ganeshie8):

when u substitute, u need to change both 'bounds' and 'differential'

ganeshie8 (ganeshie8):

just u need to practice more... i see u have the concept already :)

OpenStudy (anonymous):

Umm not sur finally

ganeshie8 (ganeshie8):

ya

OpenStudy (anonymous):

do you put ''3'' in the equation - (4 in the equation) ?

ganeshie8 (ganeshie8):

do u have skype ?

OpenStudy (anonymous):

no :(

ganeshie8 (ganeshie8):

np, i dint get ur q.... can u elaborate a bit ha :)

OpenStudy (anonymous):

\[\int\limits_{a}^{b} f(x)dx = F(b) - F(a)\]

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

so you do this with the new bounds?

ganeshie8 (ganeshie8):

yup, lets do it step by step again ok ?

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_0^1 \frac{t}{\sqrt{4-t^2}}dt}\)

ganeshie8 (ganeshie8):

ur original question is like that right ?

OpenStudy (anonymous):

Ok but when you pu the 3 in the equation, the interior of the radical is negative :(

OpenStudy (anonymous):

Yes!

ganeshie8 (ganeshie8):

thats a good observation actually ! lets see whats actually happening while solving it :)

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_0^1 \frac{t}{\sqrt{4-t^2}}dt}\) substitute \(\large u = 4-t^2\)

ganeshie8 (ganeshie8):

\(\large u = 4-t^2\) differentiate both sides

ganeshie8 (ganeshie8):

wat do u get ?

OpenStudy (anonymous):

-du/2 = t dt

ganeshie8 (ganeshie8):

yes, so we can replace tdt wid -du/2 and 4-t^2 wid u

ganeshie8 (ganeshie8):

lets do that first, we can wry about bounds afterwards :)

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_0^1 \frac{t}{\sqrt{4-t^2}}dt}\) substitute \(\large u = 4-t^2\) \(\large du = -2t dt\) \(\large -du/2 = tdt\) so the itegrand becomes : \(\large \int_{...}^{....} \frac{1}{\sqrt{u}} (-du/2)\)

ganeshie8 (ganeshie8):

let me knw if ur fine so far

OpenStudy (anonymous):

ok thanks :)

ganeshie8 (ganeshie8):

next, setup the bounds

ganeshie8 (ganeshie8):

u = 4-t^2 when t=0, u = ? when t = 1, u = ?

OpenStudy (anonymous):

4 and 3

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_0^1 \frac{t}{\sqrt{4-t^2}}dt}\) substitute \(\large u = 4-t^2\) \(\large du = -2t dt\) \(\large -du/2 = tdt\) so the itegrand becomes : \(\large \int_{...}^{....} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \int_{4}^{3} \frac{1}{\sqrt{u}} (-du/2)\)

ganeshie8 (ganeshie8):

evaluate the new integral now

ganeshie8 (ganeshie8):

factor out the constant -1/2

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_0^1 \frac{t}{\sqrt{4-t^2}}dt}\) substitute \(\large u = 4-t^2\) \(\large du = -2t dt\) \(\large -du/2 = tdt\) so the itegrand becomes : \(\large \int_{...}^{....} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \int_{4}^{3} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \frac{-1}{2}\int_{4}^{3} \frac{1}{\sqrt{u}} du\)

ganeshie8 (ganeshie8):

fine so far ? can u do the rest ha

OpenStudy (anonymous):

you + 1 the exposant of the u ( that you put in the nominator) / the answer of the addition? And after you do the thing with the soustraction ?

ganeshie8 (ganeshie8):

yes, here is the formula : \(\large \int x^n dx = \frac{x^{n+1}}{n+1}\)

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_0^1 \frac{t}{\sqrt{4-t^2}}dt}\) substitute \(\large u = 4-t^2\) \(\large du = -2t dt\) \(\large -du/2 = tdt\) so the itegrand becomes : \(\large \int_{...}^{....} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \int_{4}^{3} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \frac{-1}{2}\int_{4}^{3} \frac{1}{\sqrt{u}} du)\) \(\large \frac{-1}{2}\int_{4}^{3} u^{\frac{-1}{2} } du\)

OpenStudy (anonymous):

ok thnak you very mucch!! I'M okay with the rest :)

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_0^1 \frac{t}{\sqrt{4-t^2}}dt}\) substitute \(\large u = 4-t^2\) \(\large du = -2t dt\) \(\large -du/2 = tdt\) so the itegrand becomes : \(\large \int_{...}^{....} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \int_{4}^{3} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \frac{-1}{2}\int_{4}^{3} \frac{1}{\sqrt{u}} du)\) \(\large \frac{-1}{2}\int_{4}^{3} u^{\frac{-1}{2} } du\) \(\large \frac{-1}{2} \frac{u^{\frac{-1}{2}+1 }}{\frac{-1}{2}+1} \Big|_4^3\)

ganeshie8 (ganeshie8):

good luck wid the rest :)

OpenStudy (anonymous):

doesnt workkkkk uhuhuhh

OpenStudy (anonymous):

it gaves me : -1 - 0

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_0^1 \frac{t}{\sqrt{4-t^2}}dt}\) substitute \(\large u = 4-t^2\) \(\large du = -2t dt\) \(\large -du/2 = tdt\) so the itegrand becomes : \(\large \int_{...}^{....} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \int_{4}^{3} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \frac{-1}{2}\int_{4}^{3} \frac{1}{\sqrt{u}} du)\) \(\large \frac{-1}{2}\int_{4}^{3} u^{\frac{-1}{2} } du\) \(\large \frac{-1}{2} \frac{u^{\frac{-1}{2}+1 }}{\frac{-1}{2}+1} \Big|_4^3 \) \(\large \frac{-1}{2} \frac{u^{\frac{1}{2} }}{\frac{1}{2}} \Big|_4^3 \)

ganeshie8 (ganeshie8):

\(\large \mathbb{\int_0^1 \frac{t}{\sqrt{4-t^2}}dt}\) substitute \(\large u = 4-t^2\) \(\large du = -2t dt\) \(\large -du/2 = tdt\) so the itegrand becomes : \(\large \int_{...}^{....} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \int_{4}^{3} \frac{1}{\sqrt{u}} (-du/2)\) \(\large \frac{-1}{2}\int_{4}^{3} \frac{1}{\sqrt{u}} du)\) \(\large \frac{-1}{2}\int_{4}^{3} u^{\frac{-1}{2} } du\) \(\large \frac{-1}{2} \frac{u^{\frac{-1}{2}+1 }}{\frac{-1}{2}+1} \Big|_4^3 \) \(\large \frac{-1}{2} \frac{u^{\frac{1}{2} }}{\frac{1}{2}} \Big|_4^3 \) \(\large - \sqrt{u} \Big|_4^3 \)

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (anonymous):

\[\frac{ -1 }{ 2 } \frac{ u^{1/2} }{ 1/2

OpenStudy (anonymous):

oops but yes thats what I tap

OpenStudy (anonymous):

wrote/

ganeshie8 (ganeshie8):

yahh :)

ganeshie8 (ganeshie8):

\(\large - \sqrt{u} \Big|_4^3 \) u just need to evaluate this

OpenStudy (anonymous):

4-9= -5......

ganeshie8 (ganeshie8):

\(\large - \sqrt{u} \Big|_4^3 \) \(\large - [\sqrt{3} - \sqrt{4}] \)

ganeshie8 (ganeshie8):

\(\large - \sqrt{u} \Big|_4^3 \) \(\large - [\sqrt{3} - \sqrt{4}] \) \(\large - [\sqrt{3} - 2] \)

ganeshie8 (ganeshie8):

use ur calculator :P

OpenStudy (anonymous):

u = 4-t^2 ?

ganeshie8 (ganeshie8):

we can forget about u, cuz we changed the bounds also

ganeshie8 (ganeshie8):

\(\large - \sqrt{u} \Big|_4^3 \) \(\large - [\sqrt{3} - \sqrt{4}] \) \(\large - [\sqrt{3} - 2] \) \(\large 2 - \sqrt{3}\)

ganeshie8 (ganeshie8):

thats the final answer

OpenStudy (anonymous):

umm okay

OpenStudy (anonymous):

were suppose to change the bounds everytimes?

ganeshie8 (ganeshie8):

yes, everytime we do substitution, we need to change bounds also

ganeshie8 (ganeshie8):

im going for dinner... brb u go ahead wid other problems.. .the more u do these, the more u wil feel confident :)

OpenStudy (anonymous):

Okay thank you :p didn't know I was this much noob in those integrals..... haha i'm shy!! good dinner :)

OpenStudy (anonymous):

got problems with another question ... can you help me :)

OpenStudy (anonymous):

I can send you a picture of what I did